OpenStudy (anonymous):

Find the solution to the system of equations. 4x-3y-z=3 -3x+5y+z=9 x+2z=6

OpenStudy (anonymous):

Have you learned "elimination?"

OpenStudy (anonymous):

Yes, but I don't remember how to do it.

OpenStudy (anonymous):

Multiply the first equation by 2, which means multiply every term by 2: 4x-3y-z=3 --> 8x-6y-2z=6 do that because now you have two equations, one with +2z and one with -2z

OpenStudy (anonymous):

combine those two: 8x-6y-2z=6 (was 4x-3y-z=3) and x+2z=6 by adding like terms: 8x+x, -6y+0, -2z + 2z, 6 + 6 to get: 9x - 6y = 12

OpenStudy (anonymous):

Now you need another equation with just x, y and a constant

OpenStudy (anonymous):

4x-3y-z=3 -3x+5y+z=9 can be combined similarly, can you do it?

OpenStudy (anonymous):

Well, it's x+2y = 12

OpenStudy (anonymous):

so now you combine x+2y = 12 and 9x - 6y = 12 by multiplying x+2y = 12 by 3: 3x+6y = 36 and 9x - 6y = 12 add each term: 12x = 48 x=4 then solve for y by substitution in 9x - 6y = 12 and z by substitution in 4x-3y-z=3